# The left-to-right minima basis of the group algebra of the symmetric group (updated version)

### v1

@@ 1 Introduction
The span of the elements $\mathbf{B}_{I}$ is the classical descent algebra $\Sigma_{n}\subseteq\mathcal{A}$ introduced by Solomon in [Sol76]. We call the family $(\mathbf{B}_{I})_{I\subseteq[n-1]}$ the $\mathbf{B}$-basis of $\Sigma_{n}$. The aim of this paper is to construct and study a new basis $(\beta_{w})_{w\in S_{n}}$ of the whole algebra $\mathcal{A}$ which is compatible with the right-action of the descent algebra in the following sense: there should exist a partial order $\preceq$ on $S_{n}$ such that for every $w\in S_{n}$ and every $I\subseteq[n-1]$ one has
@@
In other words, the action of $\Sigma_{n}$ on $\mathcal{A}$ should be represented by triangular matrices on this matrix. Our basis will furthermore be compatible with a natural family of right ideals of $\mathcal{A}$ generated by the $\mathbf{B}$-basis. In the sequel we introduce a poset-indexed filtration of $\mathcal{A}$ built from the right ideals $\mathbf{B}_{\alpha}\mathcal{A}$ (reindexing $\mathbf{B}_{I}$ by compositions), construct our candidate basis, and prove that it is a basis adapted to the filtration.
The paper is organised as follows. In Section 2 we define the filtration, recall the necessary combinatorics of integer compositions, and record basic properties of the $\mathbf{B}$-basis action. Section 3 introduces the left-to-right-minima basis and states the main theorem connecting it with the filtration. The remaining sections are devoted to the proof and to examples.
@@ 2.1 Compositions and refinement
Both relations $\preceq$ and $\preceq_{\pi}$ are weak partial orders; their strict versions will be called $\prec$ and $\prec_{\pi}$.
@@ 2.2 Action of the $\mathbf{B}$-basis and right ideals
Elements of the $\mathbf{B}$-basis can be conveniently multiplied using the Solomon’s Mackey formula (see e.g. [Willigenburg98, Theorem 2]). If $M$ is a nonnegative integer matrix (i.e., matrix with nonnegative integer entries), denote by $\operatorname{rows}(M)$ and $\operatorname{cols}(M)$ its row- and column-sum vectors and by $\operatorname{read}(M)$ the composition obtained by concatenating the rows of $M$ (omitting zeros). Given $\alpha,\beta\in\operatorname{Comp}_{n}$, write $\mathbb{N}^{\alpha,\beta}$ for the set of nonnegative integer matrices with column sums $\alpha$ and row sums $\beta$. Solomon’s formula (in our notation) reads
@@
be the list of the $\gamma$-blocks. Let $\mathbf{B}_{\gamma\rightarrow\beta}$ be the sum of those permutations $w\in S_{n}$ that preserve the $\gamma$-blocks (i.e., satisfy $w\left(I_{p}\right)=I_{p}$ for all $p$) and satisfy $\operatorname*{Des}w\subseteq\operatorname*{Set}\left(\beta\right)$. By the definition of $\mathbf{B}_{\beta}$, we have
@@
where the sum runs over compositions $\alpha$ strictly finer than $\beta$, or, equivalently, over compositions $\alpha$ satisfying $\widetilde{\alpha}\prec_{\pi}\widetilde{\beta}$ (the latter sum has more addends than the former, but Lemma 1 shows that these extra summands merely duplicate existing summands). Left multiplication by $\mathbf{B}_{\alpha}$ acts on these quotient spaces by scalar multiplication:
@@
Fix $\alpha,\beta\in\operatorname{Comp}_{n}$ and let $\eta_{\beta}(\alpha)$ be the number of functions $f:[\ell(\beta)]\to[\ell(\alpha)]$ such that for every $j\in[\ell(\alpha)]$ one has $\alpha_{j}=\sum_{i\in f^{-1}(j)}\beta_{i}$. Then left multiplication by $\mathbf{B}_{\alpha}$ acts on the quotient $\mathcal{R}_{\beta}/\sum_{\alpha^{\prime}\prec\beta}\mathcal{R}_{\alpha^{\prime}}$ as multiplication by the scalar $\eta_{\beta}(\alpha)$.
@@
This is an easy corollary of Solomon’s formula (2): the matrices counted by $\eta_{\beta}(\alpha)$ are exactly those contributing the diagonal term $\mathbf{B}_{\beta}$ in $\mathbf{B}_{\alpha}\mathbf{B}_{\beta}$; hence $\mathbf{B}_{\alpha}\mathbf{B}_{\beta}=\eta_{\beta}(\alpha)\mathbf{B}_{\beta}+\left(\text{a sum of }\mathbf{B}_{\alpha^{\prime}}\text{ with }\alpha^{\prime}\prec\beta\right)$. Since $\mathcal{R}_{\beta}=\mathbf{B}_{\beta}\mathcal{A}$, this proves the claim. ∎
@@ 3.2 A new basis of the group algebra
We shall prove that these vectors form a basis of $\mathcal{A}$. The key property is the following triangularity statement.
@@
Let $\leq_{\mathrm{lex}}$ denote the lexicographic order relation on permutations seen as their one-line notation. Let $w,u\in S_{n}$ be such that $\operatorname{Des}(u)\subseteq\operatorname{LRM}^{\prime}(w)$. Then, $uw\;\leq_{\mathrm{lex}}\;w$ .
@@
Assume $u\neq\operatorname{id}$ (otherwise the claim is trivial). Choose the smallest $k$ with $u(k)<k$ (this exists since $u\neq\operatorname{id}$). Then, $k-1\in\operatorname{Des}(u)\subseteq\operatorname{LRM}^{\prime}(w)$, so that $k\in\operatorname{LRM}(w)$.
Let $\ell$ be the largest element of $\operatorname{LRM}(w)$ such that $u(\ell)<\ell$. We show that
$u(i)\leq i\qquad\text{for all }i>\ell.$ (3)
Suppose on the contrary that $u(i)>i$ for some $i>\ell$ and choose such $i$ maximal. Then all $s>i$ satisfy $u(s)\leq s$. Among these $s>i$ there must exist one with $u(s)<s$; choose the smallest such $s$. Thus, $s>i$ and $u\left(s\right)<s$ and $u\left(s-1\right)\geq s-1$ (this follows from the minimality of $s$ if s1>is-1>i, and otherwise follows from $u\left(i\right)>i$). Now, $u\left(s-1\right)\geq s-1\geq u\left(s\right)$ (since $u\left(s\right)<s$) and thus u(s1)>u(s)u(s-1)>u(s), so $s-1\in\operatorname{Des}(u)\subseteq\operatorname{LRM}^{\prime}(w)$, hence $s\in\operatorname{LRM}(w)$. This contradicts the maximality of $\ell$. Thus (3) holds.
Because $\ell\in\operatorname{LRM}(w)$, every entry appearing to the left of $\ell$ in the one-line notation of $w$ is $\ell$. Such entries $i>\ell$ satisfy $u(i)\leq i$ by (3). Combining these inequalities with $u(\ell)<\ell$ shows that $uw<_{\mathrm{lex}}w$ lexicographically. ∎
@@ 3.3 Connection with the $\mathbf{B}$-basis filtration
The mere existence of such a basis when $\mathbf{k}$ is a field is not new, and is implicit in the works of Bidigare, Brown and others. The formula for the eigenvalues can be pieced together from [RSW2014, Proposition 4.2] and [Bidigare-thesis, Corollary 4.1.3] and Foata’s fundamental transformation. Our innovation is constructing such a basis explicitly, for all commutative rings $\mathbf{k}$, and obtain a wholly new proof of Corollary 3.
@@ 4.1 Definitions
Dynkin elements are fundamental to the study of the free Lie algebra; see, e.g., [Reutenauer1993, (8.4.2)]. We shall have more use for their products in $F_{n}$.
@@ 4.2 Action on the $\mathbf{B}$-basis
Theorem 3 (Lemma 9.33 of [Reutenauer1993]).
@@
Reutenauer’s [Reutenauer1993, Lemma 9.33] is even more general: He views $F_{n}$ as the tensor algebra $T\left(V\right)$ of the free $\mathbf{k}$-module $V=\mathbf{k}^{n}$ with basis $\left(\underline{1},\underline{2},\ldots,\underline{n}\right)$. This equips $F_{n}$ with a Hopf algebra structure: that of the tensor Hopf algebra $T\left(V\right)$. All Dynkin elements $V^{S}$ are primitive elements of $T\left(V\right)$, as can be easily proved by induction on $\left|S\right|$.
@@ 5 Proof of the main theorem
Then, $\chi=\operatorname*{id}$ and $v=\operatorname*{id}$ and $\gamma=\beta$.
@@
However, our induction hypothesis yields $\beta_{j}=\gamma_{j}$ for all $j>i$, and thus easily $\beta_{\leq i}=\gamma_{\leq i}$ (since $\left|\beta\right|=n=\left|\gamma\right|$). Furthermore, it shows that $\left.v\mid_{\operatorname*{Set}\left(\beta\right)_{j}}\right.=\operatorname*{id}$ for all $j>i$, and thus $\left.v\mid_{\left[n\right]-\left[\beta_{\leq i}\right]}\right.=\operatorname*{id}$, so that $v\left(\left[n\right]-\left[\beta_{\leq i}\right]\right)=\left[n\right]-\left[\beta_{\leq i}\right]$. By taking complements, we obtain $v\left(\left[\beta_{\leq i}\right]\right)=\left[\beta_{\leq i}\right]$, thus $v^{-1}\left(\left[\beta_{\leq i}\right]\right)=\left[\beta_{\leq i}\right]$. But $vu=w$, so that $w^{-1}=u^{-1}v^{-1}$ and therefore $w^{-1}\left(\left[\beta_{\leq i}\right]\right)=u^{-1}\left(v^{-1}\left(\left[\beta_{\leq i}\right]\right)\right)=u^{-1}\left(\left[\beta_{\leq i}\right]\right)$, because $v^{-1}\left(\left[\beta_{\leq i}\right]\right)=\left[\beta_{\leq i}\right]$.
@@
In other words, $\left.v\mid_{\operatorname*{Set}\left(\gamma\right)_{i}}\right.=\operatorname*{id}$. That is, $\left.v\mid_{\operatorname*{Set}\left(\beta\right)_{i}}\right.=\operatorname*{id}$ (since $\operatorname*{Set}\left(\beta\right)_{i}=\operatorname*{Set}\left(\gamma\right)_{i}$). We have now proved all four parts of our claim $\mathcal{C}\left(i\right)$. Thus, the induction step is complete, and as we said, this finishes the proof of the lemma. ∎
@@
$\left[\underline{w}\right]\left(V^{\operatorname{Set}(\beta)_{\chi(1)}}\cdots V^{\operatorname{Set}(\beta)_{\chi(p)}}w\right)=\begin{cases}1,&\text{if }\chi=\operatorname*{id};\\ 0,&\text{if }\chi\neq\operatorname*{id},\end{cases}$ (15)
because Lemma 5 (applied to $u=w$ and $\gamma=\beta$) shows that a term $\underline{v}$ in the expansion of the product $V^{\operatorname{Set}(\beta)_{\chi(1)}}\cdots V^{\operatorname{Set}(\beta)_{\chi(p)}}$ (where $v\in S_{n}$; keep in mind that this product belongs to the image of $\mathbf{k}[S_{n}]$ under $a\mapsto\underline{a}$) can satisfy $vw=w$ (that is, $\underline{v}w=\underline{w}$) only if $\chi=\operatorname*{id}$ and $v=\operatorname*{id}$, and it is easy to see (using Proposition 3) that the coefficient of $\underline{\operatorname*{id}}$ in $V^{\operatorname{Set}(\beta)_{\chi(1)}}\cdots V^{\operatorname{Set}(\beta)_{\chi(p)}}$ is $1$ for $\chi=\operatorname*{id}$. Using (15), we can rewrite (14) as
$\left[\underline{w}\right]\left(\mathbf{V}_{\beta}\mathbf{B}_{\beta}w\right)=\sum_{\begin{subarray}{c}\chi\in S_{p};\\ \beta_{\chi\left(i\right)}=\beta_{i}\text{ for each }i\end{subarray}}\begin{cases}1,&\text{if }\chi=\operatorname*{id};\\ 0,&\text{if }\chi\neq\operatorname*{id}\end{cases}\ \ =1.$
@@
But each addend $\left[\underline{w}\right]\left(V^{\operatorname{Set}(\beta)_{\chi(1)}}\cdots V^{\operatorname{Set}(\beta)_{\chi(p)}}\cdot u\right)$ in this sum is 0, since Lemma 5 shows that a term $\underline{v}$ in the expansion of $V^{\operatorname{Set}(\beta)_{\chi(1)}}\cdots V^{\operatorname{Set}(\beta)_{\chi(p)}}$ can satisfy $vu=w$ (that is, $\underline{v}u=\underline{w}$) only if $\chi=\operatorname*{id}$ and $v=\operatorname*{id}$, but this is impossible (since $v=\operatorname*{id}$ would simplify $vu=w$ to $u=w$, contradicting our assumption $u\neq w$). Hence, the whole sum in (16) is 0, and we obtain $\left[\underline{w}\right]\left(\mathbf{V}_{\beta}\mathbf{B}_{\gamma}u\right)=0$. This proves part (b). ∎
@@
We showed that $\mathcal{S}_{\alpha}\subseteq\mathcal{R}_{\alpha}$ in Section 3.3. Thus, it remains to prove that $\mathcal{R}_{\alpha}\subseteq\mathcal{S}_{\alpha}$. So let $b\in\mathcal{R}_{\alpha}$ be arbitrary. We must show that $b\in\mathcal{S}_{\alpha}$. The family $\{\mathbf{B}_{\operatorname*{cLRM}\nolimits^{\prime}(w)}\,w\}_{w\in S_{n}}=\{\mathbf{B}_{\operatorname*{LRM}\nolimits^{\prime}(w)}\,w\}_{w\in S_{n}}$ is a basis of $\mathcal{A}$ by Corollary 2. Hence, we can expand $b$ as
@@
$=\sum_{u\in S_{n}}\lambda_{u}\,\mathbf{V}_{\beta}\mathbf{B}_{\operatorname*{cLRM}\nolimits^{\prime}(u)}\,u.$
@@
$0=\left[\underline{w}\right]\left(\sum_{u\in S_{n}}\lambda_{u}\,\mathbf{V}_{\beta}\mathbf{B}_{\operatorname*{cLRM}\nolimits^{\prime}(u)}\,u\right)=\sum_{u\in S_{n}}\lambda_{u}\,\left[\underline{w}\right]\left(\mathbf{V}_{\beta}\mathbf{B}_{\operatorname*{cLRM}\nolimits^{\prime}(u)}\,u\right).$ (18)
@@
    All the addends that satisfy $u\neq w$ but not $\mu\prec_{\pi}\widetilde{\operatorname{cLRM}^{\prime}(u)}$ are 0 (since Lemma 6 (b), applied to $\gamma=\operatorname*{cLRM}\nolimits^{\prime}\left(u\right)$, yields $\left[\underline{w}\right]\left(\mathbf{V}_{\beta}\mathbf{B}_{\operatorname*{cLRM}\nolimits^{\prime}(u)}\,u\right)=0$).

### v2

@@ 1 Introduction
The span of the elements $\mathbf{B}_{I}$ is the classical descent algebra $\Sigma_{n}\subseteq\mathcal{A}$ introduced by Solomon in [8]. We call the family $(\mathbf{B}_{I})_{I\subseteq[n-1]}$ the $\mathbf{B}$-basis of $\Sigma_{n}$. The aim of this paper is to construct and study a new basis $(\beta_{w})_{w\in S_{n}}$ of the whole algebra $\mathcal{A}$ which is compatible with the right-action of the descent algebra in the following sense: there should exist a partial order $\preceq$ on $S_{n}$ such that for every $w\in S_{n}$ and every $I\subseteq[n-1]$ one has
@@
In other words, the action of $\Sigma_{n}$ on $\mathcal{A}$ should be represented by triangular matrices on this basis. Our basis will furthermore be compatible with a natural family of right ideals of $\mathcal{A}$ generated by the $\mathbf{B}$-basis. In the sequel we introduce a poset-indexed filtration of $\mathcal{A}$ built from the right ideals $\mathbf{B}_{\alpha}\mathcal{A}$ (reindexing $\mathbf{B}_{I}$ by compositions), construct our candidate basis, and prove that it is a basis adapted to the filtration.
The paper is organised as follows. In Section 2 we define the filtration, recall the necessary combinatorics of integer compositions, and record basic properties of the $\mathbf{B}$-basis action. Section 3 introduces the left-to-right-minima basis and states the main theorem connecting it with the filtration. The remaining sections are devoted to proofs and examples.
@@ 2.1 Compositions and refinement
Both relations $\preceq$ and $\preceq_{\pi}$ are weak partial orders; their strict versions will be called $\prec$ and $\prec_{\pi}$. Note that $\alpha\preceq\beta$ implies $\widetilde{\alpha}\preceq_{\pi}\widetilde{\beta}$.
@@ 2.2 Action of the $\mathbf{B}$-basis and right ideals
Elements of the $\mathbf{B}$-basis can be conveniently multiplied using the Solomon’s Mackey formula (see e.g. [9, Theorem 2]). If $M$ is a nonnegative integer matrix (i.e., matrix with nonnegative integer entries), denote by $\operatorname{rows}(M)$ and $\operatorname{cols}(M)$ its row- and column-sum vectors and by $\operatorname{read}(M)$ the composition obtained by concatenating the rows of $M$ (omitting zeros). Given $\alpha,\beta\in\operatorname{Comp}_{n}$, write $\mathbb{N}^{\alpha,\beta}$ for the set of nonnegative integer matrices with column sums $\alpha$ and row sums $\beta$. Solomon’s formula (in our notation) reads
@@
be the list of the $\gamma$-blocks. Let $\mathbf{B}_{\gamma\rightarrow\beta}$ be the sum of those permutations $w\in S_{n}$ that preserve the $\gamma$-blocks (i.e., satisfy $w\left(I_{p}\right)=I_{p}$ for all $p$) and satisfy $\operatorname{Des}w\subseteq\operatorname*{Set}\left(\beta\right)$. By the definition of $\mathbf{B}_{\beta}$, we have
@@
where the sum runs over compositions $\alpha$ strictly finer than $\beta$, or, equivalently, over compositions $\alpha$ satisfying $\widetilde{\alpha}\prec_{\pi}\widetilde{\beta}$ (the latter sum has more addends than the former, but Lemma 1 shows that these extra addends merely duplicate existing addends). Left multiplication by $\mathbf{B}_{\alpha}$ acts on these quotient spaces by scalar multiplication:
@@
Fix $\alpha,\beta\in\operatorname{Comp}_{n}$. Let $\eta_{\beta}(\alpha)$ be the number of functions $f:[\ell(\beta)]\to[\ell(\alpha)]$ such that for every $j\in[\ell(\alpha)]$ one has $\alpha_{j}=\sum_{i\in f^{-1}(j)}\beta_{i}$. Then left multiplication by $\mathbf{B}_{\alpha}$ acts on the quotient $\mathcal{R}_{\beta}/\sum_{\alpha^{\prime}\prec\beta}\mathcal{R}_{\alpha^{\prime}}$ as multiplication by the scalar $\eta_{\beta}(\alpha)$.
@@
This is an easy corollary of Solomon’s formula (2). Indeed, all matrices $M$ on the right hand side of (2) satisfy $\operatorname{read}(M)\preceq\beta$. Among these matrices $M$, the ones that satisfy $\operatorname{read}(M)=\beta$ are precisely the ones that have exactly one nonzero entry on each row. The number of these latter matrices is $\eta_{\beta}(\alpha)$, since they can be encoded by the function $f:\left[\ell(\beta)\right]\to\left[\ell(\alpha)\right]$ where $f\left(i\right)$ is the position of the unique nonzero entry in the $i$-th row of $M$. Hence, (2) becomes
$\mathbf{B}_{\alpha}\mathbf{B}_{\beta}=\eta_{\beta}(\alpha)\mathbf{B}_{\beta}+\left(\text{a sum of }\mathbf{B}_{\alpha^{\prime}}\text{ with }\alpha^{\prime}\prec\beta\right).$ (3)
Since $\mathcal{R}_{\beta}=\mathbf{B}_{\beta}\mathcal{A}$, this proves the claim. ∎
@@ 3.2 A new basis of the group algebra
We shall prove that these vectors form a basis of $\mathcal{A}$. Two lemmas will guide our way.
@@
Let $u\in S_{n}$ be a permutation that has at least one descent. Let $m$ be the largest descent of $u$. Let $\ell=m+1$. Then:
  1. (a)
    We have $u\left(\ell\right)<\ell$.
  2. (b)
    We have $u\left(i\right)\leq i$ for all $i>\ell$.
@@
(a) Since $\ell-1=m$ is the largest descent of $u$, we have
$u\left(\ell-1\right)>u\left(\ell\right)<u\left(\ell+1\right)<u\left(\ell+2\right)<\cdots<u\left(n\right).$ (4)
Thus, the $n-\ell+1$ distinct numbers $u\left(\ell-1\right),\ u\left(\ell+1\right),\ u\left(\ell+2\right),\ \ldots,\ u\left(n\right)$ are all larger than $u\left(\ell\right)$, hence belong to the interval $\left[u\left(\ell\right)+1,\ n\right]$. Therefore, this interval must contain at least $n-\ell+1$ numbers. In other words, $\left|\left[u\left(\ell\right)+1,\ n\right]\right|\geq n-\ell+1>n$-$\ell$. Since $\left|\left[u\left(\ell\right)+1,\ n\right]\right|=n-u\left(\ell\right)$, this rewrites as $n-u\left(\ell\right)>n$-$\ell$, whence $u(\ell)<\ell$. This proves part (a).
(b) Let $i>\ell$. Then, (4) yields $u(i)<u(i+1)<\cdots<u(n)$. Thus, the nin-i distinct numbers $u\left(i+1\right),\ u\left(i+2\right),\ \ldots,\ u\left(n\right)$ are all larger than $u\left(i\right)$, hence belong to the interval $\left[u\left(i\right)+1,\ n\right]$. Therefore, this interval must contain at least nin-i numbers. In other words, $\left|\left[u\left(i\right)+1,\ n\right]\right|\geq n-i$. Since $\left|\left[u\left(i\right)+1,\ n\right]\right|=n-u\left(i\right)$, this rewrites as $n-u\left(i\right)\geq n-i$, whence $u(i)<i$. This proves part (b). ∎
Lemma 4.
Let $\leq_{\operatorname{lex}}$ denote the lexicographic order relation on permutations seen as their one-line notation. Let $w,u\in S_{n}$ be such that $\operatorname{Des}(u)\subseteq\operatorname{LRM}^{\prime}(w)$. Then, $uw\;\leq_{\operatorname{lex}}\;w$ .
Proof.
Assume that $u$ has at least one descent (since otherwise, we have $u\left(1\right)\leq u\left(2\right)\leq\cdots\leq u\left(n\right)$, therefore $u=\operatorname{id}$ and thus $uw=w$, rendering the lemma obvious). Let $m$ be the largest descent of $u$. Thus, $m\in\operatorname{Des}(u)$. Let $\ell=m+1$. Then, Lemma 3 (a) yields $u(\ell)<\ell$. Furthermore, $\ell-1=m\in\operatorname{Des}(u)\subseteq\operatorname{LRM}^{\prime}(w)$, hence $\ell\in\operatorname{LRM}(w)$. Therefore, every entry $i$ appearing to the left of $\ell$ in the one-line notation of $w$ is $\ell$, and thus satisfies $u(i)\leq i$ by Lemma 3 (b). Combining this with $u(\ell)<\ell$, we see that $uw<_{\operatorname{lex}}w$ lexicographically. This proves Lemma 4. ∎
@@
Example 2.
Let us write permutations $\sigma\in S_{n}$ in one-line notation, i.e., as $\left[\sigma\left(1\right)\ \sigma\left(2\right)\ \cdots\ \sigma\left(n\right)\right]$. For $n=3$, the basis $\bigl\{\mathbf{B}_{\operatorname{LRM}^{\prime}(w)}\,w\;\big|\;w\in S_{n}\bigr\}$ of $\mathcal{A}$ consists of the vectors
$\mathbf{B}_{\operatorname{LRM}^{\prime}(\left[123\right])}\,\left[123\right]$ $=\mathbf{B}_{\left(3\right)}\ \left[123\right]=\left[123\right];$
$\mathbf{B}_{\operatorname{LRM}^{\prime}(\left[132\right])}\,\left[132\right]$ $=\mathbf{B}_{\left(3\right)}\ \left[132\right]=\left[132\right];$
$\mathbf{B}_{\operatorname{LRM}^{\prime}(\left[213\right])}\,\left[213\right]$ $=\mathbf{B}_{\left(1,2\right)}\ \left[213\right]=\left[213\right]+\left[123\right]+\left[132\right];$
$\mathbf{B}_{\operatorname{LRM}^{\prime}(\left[231\right])}\,\left[231\right]$ $=\mathbf{B}_{\left(1,2\right)}\ \left[231\right]=\left[231\right]+\left[132\right]+\left[123\right];$
$\mathbf{B}_{\operatorname{LRM}^{\prime}(\left[312\right])}\,\left[312\right]$ $=\mathbf{B}_{\left(2,1\right)}\ \left[312\right]=\left[312\right]+\left[213\right]+\left[123\right];$
$\mathbf{B}_{\operatorname{LRM}^{\prime}(\left[321\right])}\,\left[321\right]$ $=\mathbf{B}_{\left(1,1,1\right)}\ \left[321\right]=\left[321\right]+\left[312\right]+\left[231\right]+\left[213\right]+\left[132\right]+\left[123\right].$
@@ 3.3 Connection with the $\mathbf{B}$-basis filtration
(This is an outline; see Section 6 for a more detailed statement.)
@@
The mere existence of such a basis when $\mathbf{k}$ is a field is not new, and is implicit in the works of Bidigare, Brown and others. The formula for the eigenvalues can be pieced together from [5, Proposition 4.2] and [1, Corollary 4.1.3] and Foata’s fundamental transformation. Our innovation is constructing such a basis explicitly, for all commutative rings $\mathbf{k}$; this yields a wholly new proof of Corollary 3.
@@ 4.1 Definitions
Dynkin elements are fundamental to the study of the free Lie algebra; see, e.g., [6, (8.4.2)]. We shall have more use for their products in $F_{n}$.
@@ 4.2 Action on the $\mathbf{B}$-basis
Theorem 3 (Lemma 9.33 of [6], or Theorem 2.1 of [3]).
@@
Reutenauer’s [6, Lemma 9.33] is even more general: He views $F_{n}$ as the tensor algebra $T\left(V\right)$ of the free $\mathbf{k}$-module $V=\mathbf{k}^{n}$ with basis $\left(\underline{1},\underline{2},\ldots,\underline{n}\right)$. This equips $F_{n}$ with a Hopf algebra structure: that of the tensor Hopf algebra $T\left(V\right)$. All Dynkin elements $V^{S}$ are primitive elements of $T\left(V\right)$, as can be easily proved by induction on $\left|S\right|$.
@@ 5 Proof of the main theorem
Then, $\chi=\operatorname{id}$ and $v=\operatorname{id}$ and $\gamma=\beta$.
@@
However, our induction hypothesis yields $\beta_{j}=\gamma_{j}$ for all $j>i$, and thus easily $\beta_{\leq i}=\gamma_{\leq i}$ (since $\left|\beta\right|=n=\left|\gamma\right|$). Furthermore, it shows that $\left.v\mid_{\operatorname*{Set}\left(\beta\right)_{j}}\right.=\operatorname{id}$ for all $j>i$, and thus $\left.v\mid_{\left[n\right]-\left[\beta_{\leq i}\right]}\right.=\operatorname{id}$, so that $v\left(\left[n\right]-\left[\beta_{\leq i}\right]\right)=\left[n\right]-\left[\beta_{\leq i}\right]$. By taking complements, we obtain $v\left(\left[\beta_{\leq i}\right]\right)=\left[\beta_{\leq i}\right]$, thus $v^{-1}\left(\left[\beta_{\leq i}\right]\right)=\left[\beta_{\leq i}\right]$. But $vu=w$, so that $w^{-1}=u^{-1}v^{-1}$ and therefore $w^{-1}\left(\left[\beta_{\leq i}\right]\right)=u^{-1}\left(v^{-1}\left(\left[\beta_{\leq i}\right]\right)\right)=u^{-1}\left(\left[\beta_{\leq i}\right]\right)$, because $v^{-1}\left(\left[\beta_{\leq i}\right]\right)=\left[\beta_{\leq i}\right]$.
@@
In other words, $\left.v\mid_{\operatorname*{Set}\left(\gamma\right)_{i}}\right.=\operatorname{id}$. That is, $\left.v\mid_{\operatorname*{Set}\left(\beta\right)_{i}}\right.=\operatorname{id}$ (since $\operatorname*{Set}\left(\beta\right)_{i}=\operatorname*{Set}\left(\gamma\right)_{i}$). We have now proved all four parts of our claim $\mathcal{C}\left(i\right)$. Thus, the induction step is complete, and as we said, this finishes the proof of the lemma. ∎
@@
$\left[\underline{w}\right]\left(V^{\operatorname{Set}(\beta)_{\chi(1)}}\cdots V^{\operatorname{Set}(\beta)_{\chi(p)}}w\right)=\begin{cases}1,&\text{if }\chi=\operatorname{id};\\ 0,&\text{if }\chi\neq\operatorname{id},\end{cases}$ (16)
because Lemma 6 (applied to $u=w$ and $\gamma=\beta$) shows that a term $\underline{v}$ in the expansion of the product $V^{\operatorname{Set}(\beta)_{\chi(1)}}\cdots V^{\operatorname{Set}(\beta)_{\chi(p)}}$ (where $v\in S_{n}$; keep in mind that this product belongs to the image of $\mathbf{k}[S_{n}]$ under $a\mapsto\underline{a}$) can satisfy $vw=w$ (that is, $\underline{v}w=\underline{w}$) only if $\chi=\operatorname{id}$ and $v=\operatorname{id}$, and it is easy to see (using Proposition 3) that the coefficient of $\underline{\operatorname{id}}$ in $V^{\operatorname{Set}(\beta)_{\chi(1)}}\cdots V^{\operatorname{Set}(\beta)_{\chi(p)}}$ is $1$ for $\chi=\operatorname{id}$. Using (16), we can rewrite (15) as
$\left[\underline{w}\right]\left(\mathbf{V}_{\beta}\mathbf{B}_{\beta}w\right)=\sum_{\begin{subarray}{c}\chi\in S_{p};\\ \beta_{\chi\left(i\right)}=\beta_{i}\text{ for each }i\end{subarray}}\begin{cases}1,&\text{if }\chi=\operatorname{id};\\ 0,&\text{if }\chi\neq\operatorname{id}\end{cases}\ \ =1.$
@@
But each addend $\left[\underline{w}\right]\left(V^{\operatorname{Set}(\beta)_{\chi(1)}}\cdots V^{\operatorname{Set}(\beta)_{\chi(p)}}\cdot u\right)$ in this sum is 0, since Lemma 6 shows that a term $\underline{v}$ in the expansion of $V^{\operatorname{Set}(\beta)_{\chi(1)}}\cdots V^{\operatorname{Set}(\beta)_{\chi(p)}}$ can satisfy $vu=w$ (that is, $\underline{v}u=\underline{w}$) only if $\chi=\operatorname{id}$ and $v=\operatorname{id}$, but this is impossible (since $v=\operatorname{id}$ would simplify $vu=w$ to $u=w$, contradicting our assumption $u\neq w$). Hence, the whole sum in (17) is 0, and we obtain $\left[\underline{w}\right]\left(\mathbf{V}_{\beta}\mathbf{B}_{\gamma}u\right)=0$. This proves part (b). ∎
@@
We showed that $\mathcal{S}_{\alpha}\subseteq\mathcal{R}_{\alpha}$ in Section 3.3. Thus, it remains to prove that $\mathcal{R}_{\alpha}\subseteq\mathcal{S}_{\alpha}$. So let $b\in\mathcal{R}_{\alpha}$ be arbitrary. We must show that $b\in\mathcal{S}_{\alpha}$. The family $\{\mathbf{B}_{\operatorname{cLRM}^{\prime}(w)}\,w\}_{w\in S_{n}}=\{\mathbf{B}_{\operatorname*{LRM}\nolimits^{\prime}(w)}\,w\}_{w\in S_{n}}$ is a basis of $\mathcal{A}$ by Corollary 2. Hence, we can expand $b$ as
@@
$=\sum_{u\in S_{n}}\lambda_{u}\,\mathbf{V}_{\beta}\mathbf{B}_{\operatorname{cLRM}^{\prime}(u)}\,u.$
@@
$0=\left[\underline{w}\right]\left(\sum_{u\in S_{n}}\lambda_{u}\,\mathbf{V}_{\beta}\mathbf{B}_{\operatorname{cLRM}^{\prime}(u)}\,u\right)=\sum_{u\in S_{n}}\lambda_{u}\,\left[\underline{w}\right]\left(\mathbf{V}_{\beta}\mathbf{B}_{\operatorname{cLRM}^{\prime}(u)}\,u\right).$ (19)
@@
    All the addends that satisfy $u\neq w$ but not $\mu\prec_{\pi}\widetilde{\operatorname{cLRM}^{\prime}(u)}$ are 0 (since Lemma 7 (b), applied to $\gamma=\operatorname*{cLRM}\nolimits^{\prime}\left(u\right)$, yields $\left[\underline{w}\right]\left(\mathbf{V}_{\beta}\mathbf{B}_{\operatorname{cLRM}^{\prime}(u)}\,u\right)=0$).